OpenStudy (anonymous):

Okay, I have another Algebra problem that I am not sure how to solve... If someone could help me with it and help explain it to me then that would be great! :D This is the question: Which inequality models this problem? The length of a rectangle is five times its width. If the perimeter is at most 96 centimeters, what is the greatest possible value for the width? And these are the answer choices: A. 2w + 2 • (5w) ≤ 96 B. 2w + 2 • (5w) ≥ 96 C. 2w + 2 • (5w) < 96 D. 2w + 2 • (5w) > 96

OpenStudy (anonymous):

I am not sure if the answer is A. or B. What do you think?

OpenStudy (texaschic101):

P = 2L + 2W L = 5W P <= 96 2(5W) + 2W <= 96 which is the same as 2W + 2*5W <= 96 ....its A

OpenStudy (texaschic101):

when it says " at most ", it means less then or equal to

OpenStudy (anonymous):

Oh, okay now I understand! Thanks so much for you help! :D

OpenStudy (texaschic101):

anytime :)